# Fixing That Free Lunch: When, Where, and Why Synthetic Data Fails in Model-Based Policy Optimization

### v1

@@ 5.1 Ablations: Do we really need both?
To test whether the remedies identified in Sections 4.2 and 4.3 generalize beyond a single case, we performed ablations across the seven DMC tasks that MBPO previously failed in, as shown in Appendix C. FTFL succeeds in five of seven cases, enabling stable policy improvement where MBPO had previously failed. The two remaining environments, where both MBPO and STFL fail, are discussed in detail in Section 5.3. These results confirm that the failures identified earlier were not isolated artifacts but systemic issues that FTFL resolves in the majority of cases.
@@ 5.2 Generalization and Restoring the Promise of MBPO
Figure 4: Return performance on six DMC tasks. Following Agarwal et al. (2022), solid lines show interquartile mean (IQM) returns aggregated across six seeds, and shaded regions denote 95% bootstrapped confidence intervals. Results are shown for FTFL with the original model and for Tuned FTFL with increased model capacity (effective only on humanoid tasks). One result was omitted for brevity, but raw averaged returns are provided for all seven tasks in Appendix B.
Figure 5: Return performance on six OpenAI Gym tasks, using the same IQM and confidence interval conventions as Figure 4. Raw averaged returns are provided in Appendix B.
@@ 5.3 Remaining Failures and Towards a Taxonomy of RL Failure Modes
While FTFL resolves the failure modes in quadruped and humanoid environments, it does not fully mitigate MBPO’s issues on hopper-hop and hopper-stand. In these Hopper tasks, scale and residual issues are corrected, reward collapse is prevented, and variance is reduced, consistent with the results in Section 5.1. Yet performance still stalls. An additional experiment (cf. Appendix D) shows that removing the contact sensor in the hopper observation (a discontinuous contact-force signal) allows FTFL to match SAC on hopper-stand, but not on hopper-hop, pointing to a modality-specific challenge unique to Hopper that was absent in other failing DMC environments.
@@ 6 Broader Implications and Conclusion
Usage of Large Language Models (LLMs)
@@
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our results. Implementation details for SAC, MBPO, and FTFL are provided in Appendix A, with complete hyperparameter tables included in the appendix. Additional raw return curves, remediation ablations, and environment-specific analyses are reported in Appendix B, Appendix C, and Appendix D, respectively. Our results build on the efficient MBPO reimplementation from Barkley & Fridovich-Keil (2025), which is publicly available.
Appendix A SAC and MBPO Implementation Details
@@
Appendix B Raw Return Curves for SAC, MBPO, FTFL, and Tuned FTFL in Gym and DMC
@@
Appendix C Remediation Ablation
@@
Appendix D FTFL and SAC Performance in Hopper Environments Without Touch Sensor Observations

### v2

@@ 5.1 Ablations: Do we really need both?
To test whether the remedies identified in Sections 4.2 and 4.3 generalize beyond a single case, we performed ablations across the seven DMC tasks that MBPO previously failed in, as shown in Appendix E. FTFL succeeds in five of seven cases, enabling stable policy improvement where MBPO had previously failed. The two remaining environments, where both MBPO and STFL fail, are discussed in detail in Section 5.3. These results confirm that the failures identified earlier were not isolated artifacts but systemic issues that FTFL resolves in the majority of cases.
@@ 5.2 Generalization and Restoring the Promise of MBPO
Figure 4: Return performance on six DMC tasks. Following Agarwal et al. (2022), solid lines show interquartile mean (IQM) returns aggregated across six seeds, and shaded regions denote 95% bootstrapped confidence intervals. Results are shown for FTFL with the original model and for Tuned FTFL with increased model capacity (effective only on humanoid tasks). One result was omitted for brevity, but raw averaged returns are provided for all seven tasks in Appendix D.
Figure 5: Return performance on six OpenAI Gym tasks, using the same IQM and confidence interval conventions as Figure 4. Raw averaged returns are provided in Appendix D.
@@ 5.3 Remaining Failures and Towards a Taxonomy of RL Failure Modes
While FTFL resolves the failure modes in quadruped and humanoid environments, it does not fully mitigate MBPO’s issues on hopper-hop and hopper-stand. In these Hopper tasks, scale and residual issues are corrected, reward collapse is prevented, and variance is reduced, consistent with the results in Section 5.1. Yet performance still stalls. An additional experiment (cf. Appendix F) shows that removing the contact sensor in the hopper observation (a discontinuous contact-force signal) allows FTFL to match SAC on hopper-stand, but not on hopper-hop, pointing to a modality-specific challenge unique to Hopper that was absent in other failing DMC environments.
@@ 6 Broader Implications and Conclusion
7 Acknowledgments
This work was supported by the National Science Foundation under Grant No. 2409535.
Appendix A Usage of Large Language Models (LLMs)
@@
Appendix B Reproducibility Statement
We have taken several steps to ensure the reproducibility of our results. Implementation details for SAC, MBPO, and FTFL are provided in Appendix C, with complete hyperparameter tables included in the appendix. Additional raw return curves, remediation ablations, and environment-specific analyses are reported in Appendix D, Appendix E, and Appendix F, respectively. Our results build on the efficient MBPO reimplementation from Barkley & Fridovich-Keil (2025), which is publicly available.
Appendix C SAC and MBPO Implementation Details
@@
Appendix D Raw Return Curves for SAC, MBPO, FTFL, and Tuned FTFL in Gym and DMC
@@
Appendix E Remediation Ablation
@@
Appendix F FTFL and SAC Performance in Hopper Environments Without Touch Sensor Observations